Golly Hospital purchased equipment with the following costs:

Purchase Date     Cost           Required Useful Life (Years)

1/1/X1                 $100,000                               8

4/1/X1                  200,000                                10

7/1/X1                 300,000                                   6

Required: All equipment has a 20 percent salvage value. What is depreciation expense for 20X1?
Answer #1

Answers

  • Equipment 1 Depreciation
    = [($100000 – 20%) / 8 years] x 12/12 = $ 10,000
  • Equipment 2 Depreciation
    = [($200000 – 20%) / 10 years] x 9 months/12 = $ 12,000
  • Equipment 3 Depreciation
    = [($300000 – 20%) / 6 years] x 6 months/12 = $ 20,000
  • Depreciation expense for 20X1
    = 10000 + 12000 + 20000
    = $ 42,000 Answer